OpenStudy (chrisga):

LaToya decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a beginner's fee of $100 and a monthly fee of $38. If x represents the number of months that LaToya is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time: 100 + 38x = C LaToya has allocated a maximum of $404 to spend on her gym membership. Which number line shows the possible number of months that LaToya can be a member of the gym?

OpenStudy (chrisga):

A number line shows numbers from 0 to 12 in increments of 1. The portion of the line from 0 to 8 is bold with shaded circles at 0 and at 8 A number line shows numbers from 0 to 13 in increments of 1. The portion of the line from 1 to 13 is bold with shaded circles at 1 and at 13 A number line shows numbers from 0 to 13 in increments of 1. The portion of the line from 0 to 13 is bold with shaded circles at 0 and at 13 A number line shows numbers from 0 to 12 in increments of 1. The portion of the line from 1 to 8 is bold with shaded circles at 1 and at 8

OpenStudy (chrisga):

I got the work but im not sure if it is A or D 100+38x=404 100-100+38x=404-100 38x=303 38/38x=304/38 x=8

OpenStudy (chrisga):

@iloveyou;*

OpenStudy (anonymous):

i suck at math lol sorry :/

OpenStudy (anonymous):

@HectorH pls help him

OpenStudy (chrisga):

can you help? @HectorH

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

your work is correct

OpenStudy (chrisga):

which one would it be A or D?

OpenStudy (anonymous):

it would be D

OpenStudy (chrisga):

OK thx alot

OpenStudy (anonymous):

no prob